# Connection between galaxy morphology and dark-matter halo structure II: predicting disk structure from dark-matter halo properties

### v1

@@ Abstract
We investigate how galactic disk structures connect to the detailed properties of their host dark-matter halos using the TNG50 simulation. From the hydrodynamic and matched dark-matter–only runs, we measure a comprehensive list of halo properties describing density structure, angular momentum, shape, assembly history, and environment. Using the morphological decomposition developed in Paper I, we quantify the sizes, scale heights, and mass fractions of the disk components for galaxies at $0\leq z\leq 4$. Random Forest (RF) regression shows halo properties alone predict disk size and thickness with high accuracy, while Symbolic Regression (SR) provides compact empirical relations with slightly lower accuracy. Disk height is consistently easier to predict than disk size, and lower-mass halos yield higher accuracy than massive halos. Predictions based on halo properties measured in the full-physics hydro simulations outperform those based on the matched dark-matter–only halos, reflecting the imprint of baryonic restructuring on the inner halo. SHAP analysis reveals that the most informative halo parameters include concentration, Einasto shape, global and inner spin, and recent mass accretion, though their importance varies across disk properties. We show that correlations between disk size and the density-profile shape arise primarily from disk-induced modification of the inner halo, rather than a primordial connection. Finally, we point out that disks become more extended with respect to their host halos at higher redshift in low-mass halos, while massive high-redshift halos show the opposite trend. We provide SR-based prescriptions that accurately map halo properties to disk structures, offering practical tools for galaxy–halo modeling.
@@ 1 Introduction
Stellar disks are particularly revealing, as their sizes (e.g. Kravtsov13; Huang17), thicknesses (e.g. Zasov02; Sotnikova06), and mass fractions (e.g. Kim13) trace the angular-momentum (AM) acquisition and dynamical heating within halos. Previous studies, including Paper I of this series (Liang25), have shown that, besides halo mass, which is the primary driver of galaxy size, secondary halo properties including spin (e.g. Fall80; Fall83; Mo98), concentration (e.g. Jiang19), and accretion rate (e.g. Dubois16) all affect disk sizes, but a comprehensive, multidimensional understanding of how halo conditions shape detailed disk structures is still lacking.
Our goals in this series of studies are to determine how well disk structure can be predicted from halo properties using cosmological simulations, to identify the most influential parameters, and to derive empirical relations that can be incorporated into semi-analytic galaxy-formation models (e.g. Cole00; Lacey16; Henriques20). To this end, we have developed an algorithm that measures the physical morphologies of simulated galaxies by decomposing them into kinematic components, and we also measure a compressive list of halo structural properties using particle data, which complements the information available from typical halo-finding algorithms (Paper I). In this work, the second paper of this series, we make use of these detailed morphological and structural measurements, and apply machine-learning techniques to uncover complex nonlinear relationships in such high-dimensional data. We adopt a framework that combines Random Forest regression (Breiman01), SHapley Additive Explanations (Lundberg17), and Symbolic Regression (e.g. Koza92) to quantify and interpret the dependence of disk morphology on halo parameters.
The paper is organized as follows. Section 2 describes the simulation data, the measurements of galaxy and halo properties, and the algorithms employed in this work. Section 3 presents the connections between galaxies and dark-matter halos extracted from our measurements, and identifies the most important halo parameters that regulate disk properties. Section 4 explores the physical reasons for the relations obtained in Section 3, and Section 5 summarizes our finding. Throughout this study, we define halos as spherical over-densities of 200 times the critical density of the Universe, and adopt the cosmological parameters assumed in the TNG simulations.
2 Method
In this section, we describe the cosmological simulation sample, define all the halo and galaxy properties considered in this study, and present the machine learning algorithms applied to these data products. For readers who wish to proceed to the results in Section 3 without delving into technical details, we provide a summary of our measurements in Table 1 and a workflow of the analysis pipeline in Fig. 1.
@@ 2 Method
2.1 Simulation and sample selection
We adopt the highest-resolution run of the Illustris-TNG suite (hereafter TNG50, Nelson19; Pillepich19), which has a gas particle mass of $10^{4.93}M_{\odot}$, a dark matter (DM) particle mass of $10^{5.65}M_{\odot}$, and a gravitational softening length of 0.288 (0.074) comoving kpc for collisionless (gas) particles. The simulation includes sub-grid models that capture the key physical processes of gas cooling, star formation, chemical enrichment, and feedback. Halos are identified using the Friends-of-Friends algorithm (FoF, Davis85) and the SUBFIND algorithm (Springel01), and are linked across snapshots via the SUBLINK merger tree algorithm (Rodriguez-Gomez15). The corresponding DM-only (DMO) simulation shares the same initial conditions, and halos are matched to their hydro-simulation counterparts using the public Subhalo Matching To Dark catalog [^1] https://www.tng-project.org/data/docs/specifications/#sec5d.
@@ 2.2 DM halo measurements
  • Halo concentration is defined as $c=R_{\rm vir}/r_{-2}$, where $r_{-2}$ is the radius where the density slope equals –2. We fit the Einasto65 profile
@@
    following Bullock01, where $j_{\rm vir}$ is the specific angular momentum within the virial radius and $V_{\rm vir}$ the circular velocity at $R_{\rm vir}$. We also measure an inner spin, $\lambda_{\rm inner}=j_{\rm vir,inner}/(\sqrt{2}R_{\rm vir}V_{\rm vir})$, using the specific angular momentum enclosed within $0.1R_{\rm vir}$, $j_{\rm vir,inner}$.
  • The mass accretion rate and mass assembly history (MAH) of a dark-matter halo are characterized using two approaches. The MAH is defined as the mass of the main branch as a function of redshift, $M_{\rm vir}(z)$. Following McBride09, we fit it with
@@
    Alternatively, the MAH can be expressed using the mass variable (vdBosch02; Wechsler02; Zhao03),
@@
    where $\sigma(M)$ is the standard deviation of the linear over-density field on mass scale $M$. Following Chen20, we apply Principal Component Analysis (PCA) to $s(z)$ at fixed redshift to reduce dimensionality. The first two principal components (PC1 and PC2) accurately capture both the shape and evolution of a typical MAH, with PC1 representing the overall monotonic growth, while PC2 encodeing a higher-order, non-monotonic deviation. Thus, we use $\dot{M}_{\rm norm}$ to trace recent accretion and (PC1, PC2) to describe the global MAH shape over cosmic time.
  • Halo formation redshift $z_{\rm 1/2}$ is defined as the highest redshift at which the main branch assembled half of its final mass $M_{\rm vir}$ (Wang11).
  • Halo 3D shape parameters are derived from the eigenvalues of the inertia tensor (Allgood06),
@@
    where the sum runs over all DM particles within the ellipsoid, and $M=\sum_{k}m_{k}$ is the enclosed mass. The eigenvalues of $\mathcal{S}$ correspond to the squared semi-axes ($a\geq b\geq c$), obtained iteratively from the virial sphere until convergence (Tomassetti16). We define the 3D axis ratios
@@
  • Environment classification is based on a halo’s location within the cosmic web. We identify web types using the eigenvalues of the deformation tensor (e.g., Hahn07; Forero-Romero09): the number of eigenvalues exceeding a threshold $\lambda_{\rm th}=0.4$ determines the environment, with 0, 1, 2, and 3 corresponding to voids, sheets, filaments, and knots, respectively. The overdensity field is computed on a $512^{3}$ grid via a cloud-in-cell interpolation, smoothed with a Gaussian filter of $R_{\rm G}=0.5\,h^{-1}{\rm ckpc}$. These parameter choices yield a visually consistent web morphology and agree with previous studies (e.g., Martizzi19).
@@ 2.3 Galaxy Quantities
2.4 Machine Learning Algorithms
@@
While (semi-)analytical and empirical models have used connections between disk size and halo spin (e.g. Cole00; Lacey16; Henriques20) or concentration (e.g. Nadler20; Behroozi22; Somerville25), it remains an open question which one is more influential (Jiang19) and whether incorporating a broader range of halo properties can enhance prediction accuracy (Liang25). Moreover, it is not yet known whether disk vertical height and mass fraction, can be robustly inferred from halo properties. Addressing these issues necessitates a systematic investigation of correlations across a large set of halo conditions, a task that can be effectively done by supervised machine-learning (ML) algorithms.
@@
In this study, we use ML algorithms to uncover relations between disk parameters and halo conditions. The targets and features are galaxy quantities and halo quantities, respectively, which are summarized in Table 1. We consider two ML algorithms. We first train the random forest (RF) regressor (Breiman01) for all halo conditions to predict our targets. Then we use SHapley Additive Explanations (SHAP) values (Lundberg17) to identify important halo properties that affect each prediction. Finally we use symbolic regression (SR, e.g. Koza92) with the most important 10 halo properties as identified by the RF regressor as input to derive empirical functions relating the targets and features. The technical details and hyper-parameters that we tune are described in Appendix A. The workflow in this work is presented in Fig. 1.
@@
For both RF and SR, we use 80% of the total samples as the training set and the remaining 20% as the test set. All features and target variables are scaled to dimensionless quantities. For the targets such as disk scale radii, disk scale heights, and galaxy sizes, they are normalized by the virial radius, and the dimensionless quantities are referred to as compactness or thickness. The stellar mass is normalized by the virial mass, while the SFR is normalized by the average halo growth rate over Hubble time. For the features, the virial mass is scaled by $10^{10}M_{\odot}$, and the other properties are already dimensionless. Except for disk mass fractions, a base-10 logarithm is applied to the target variables.
@@
3 Galaxy-halo connections
In this section, we present the models for predicting galaxy quantities with DM-halo quantities, as identified with the random-forest (RF) and symbolic-regression (SR) algorithms. We first test the performance of the regression models for global properties, specifically, stellar mass, star formation rate (SFR), and size, as a validation of the methods in Section 3.1, and then extend the analysis to disk properties in Section 3.2. For each relation, we first show the accuracy of our prediction, and then discuss the correlation between the target and the features.
3.1 Predicting galaxy global properties with halo parameters
3.1.1 Stellar mass
@@
3.1.2 Star formation rate
@@ 3.1.3 Galaxy size
Feature importance: Unlike stellar mass or SFR, the most important halo property for galaxy compactness is not virial mass but redshift. In addition, the importance of virial mass is much smaller for the hydro halos compared to the DMO halos. For the hydro halos, the role of the Einasto shape index ($\alpha$) is significant, ranking the second, whereas it is not in the top 10 in the DMO case. This difference persists in predictions for disk structures (Section 3.2), and implies that $\alpha$ is significantly influenced by the galaxy rather than being a driver of galaxy size. Other commonly important properties across both hydro halos and DMO halos are concentration $c$, spin $\lambda$, and the inner-halo spin $\lambda_{\rm inner}$. A few environmental and assembly-history properties also weakly contribute to the prediction.
@@
3.2 Predicting disk properties with halo properties
3.2.1 Disk size

### v2

@@ Abstract
We investigate how galactic disk structures connect to the detailed properties of their host dark-matter halos using the TNG50 simulation. From the hydrodynamic and matched dark-matter–only runs, we measure a comprehensive list of halo properties describing density structure, angular momentum, shape, assembly history, and environment. Using the morphological decomposition developed in Paper I, we quantify the sizes, scale heights, and mass fractions of the disk components for galaxies at $0\leq z\leq 4$. Random Forest (RF) regression shows that halo properties alone predict disk size and thickness with high accuracy, while Symbolic Regression (SR) provides compact empirical relations with slightly lower accuracy. Disk height is consistently easier to predict than disk size, and lower-mass halos yield higher accuracy than massive halos. Predictions based on halo properties measured in the hydro simulations outperform those based on halos matched in the dark-matter–only simulation, reflecting the imprint of baryonic restructuring on the inner halo. SHAP analysis reveals the most informative halo parameters include concentration, Einasto shape, global and inner spin, and recent mass accretion, though their importance varies across disk properties. We show correlations between disk size and the density-profile shape arise primarily from disk-induced modification of the inner halo, rather than a primordial connection. Finally, we point out that disks become more extended with respect to their host halos at higher redshift in low-mass halos, while massive high-redshift halos show the opposite trend. We provide SR-based prescriptions that accurately map halo properties to disk structures, offering practical tools for galaxy–halo modeling.
@@ I Introduction
Stellar disks are particularly revealing, as their sizes (e.g. Kravtsov, 2013; Huang et al., 2017), thicknesses (e.g. Zasov et al., 2002; Sotnikova and Rodionov, 2006), mass fractions (e.g. Kim and Lee, 2013; Romeo et al., 2020), and instability (e.g. Romeo et al., 2023) trace the angular-momentum (AM) acquisition, retention and dynamical heating within halos. Previous studies, including Paper I of this series (Liang et al., 2025), have shown that, besides halo mass, which is the primary driver of galaxy size, secondary halo properties including spin (e.g. Fall and Efstathiou, 1980; Fall, 1983; Mo et al., 1998), concentration (e.g. Jiang et al., 2019), and accretion rate (e.g. Dubois et al., 2016) all affect disk sizes, but a comprehensive, multidimensional understanding of how halo conditions shape detailed disk structures is still lacking.
Our goals in this series of studies are to determine how well disk structure can be predicted from halo properties using cosmological simulations, to identify the most influential parameters, and to derive empirical relations that can be incorporated into semi-analytic galaxy-formation models (e.g. Cole et al., 2000; Lacey et al., 2016; Henriques et al., 2020). To this end, we have developed an algorithm that measures the physical morphologies of simulated galaxies by decomposing them into kinematic components, and we also measure a compressive list of halo structural properties using particle data, which complements the information available from typical halo-finding algorithms (Paper I). In this work, the second paper of this series, we make use of these detailed morphological and structural measurements, and apply machine-learning techniques to uncover complex nonlinear relationships in such high-dimensional data. We adopt a framework that combines Random Forest regression (Breiman, 2001), SHapley Additive Explanations (Lundberg and Lee, 2017), and Symbolic Regression (e.g. Koza, 1992) to quantify and interpret the dependence of disk morphology on halo parameters.
The paper is organized as follows. Section II describes the simulation data, the measurements of galaxy and halo properties, and the algorithms employed in this work. Section III presents the connections between galaxies and dark-matter halos extracted from our measurements, and identifies the most important halo parameters that regulate disk properties. Section IV explores the physical reasons for the relations obtained in Section III, and Section V summarizes our finding. Throughout this study, we define halos as spherical over-densities of 200 times the critical density of the Universe, and adopt the cosmological parameters assumed in the TNG simulations.
II Method
In this section, we describe the cosmological simulation sample, define all the halo and galaxy properties considered in this study, and present the machine learning algorithms applied to these data products. For readers who wish to proceed to the results in Section III without delving into technical details, we provide a summary of our measurements in Table 1 and a workflow of the analysis pipeline in Fig. 1.
@@ II Method
II.1 Simulation and sample selection
We adopt the highest-resolution run of the Illustris-TNG suite (hereafter TNG50, Nelson et al., 2019; Pillepich et al., 2019), which has a gas particle mass of $10^{4.93}M_{\odot}$, a dark matter (DM) particle mass of $10^{5.65}M_{\odot}$, and a gravitational softening length of 0.288 (0.074) comoving kpc for collisionless (gas) particles. The simulation includes sub-grid models that capture the key physical processes of gas cooling, star formation, chemical enrichment, and feedback. Halos are identified using the Friends-of-Friends algorithm (FoF, Davis et al., 1985) and the SUBFIND algorithm (Springel et al., 2001), and are linked across snapshots via the SUBLINK merger tree algorithm (Rodriguez-Gomez et al., 2015). The corresponding DM-only (DMO) simulation shares the same initial conditions, and halos are matched to their hydro-simulation counterparts using the public Subhalo Matching To Dark catalog [^1] https://www.tng-project.org/data/docs/specifications/#sec5d.
@@ II.2 DM halo measurements
  • Halo concentration is defined as $c=R_{\rm vir}/r_{-2}$, where $r_{-2}$ is the radius where the density slope equals –2. We fit the Einasto (1965) profile
@@
    following Bullock et al. (2001), where $j_{\rm vir}$ is the specific angular momentum within the virial radius and $V_{\rm vir}$ the circular velocity at $R_{\rm vir}$. We also measure an inner spin, $\lambda_{\rm inner}=j_{\rm vir,inner}/(\sqrt{2}R_{\rm vir}V_{\rm vir})$, using the specific angular momentum enclosed within $0.1R_{\rm vir}$, $j_{\rm vir,inner}$.
  • The mass accretion rate and mass assembly history (MAH) of a dark-matter halo are characterized using two approaches. The MAH is defined as the mass of the main branch as a function of redshift, $M_{\rm vir}(z)$. Following McBride et al. (2009), we fit it with
@@
    Alternatively, the MAH can be expressed using the mass variable (van den Bosch, 2002; Wechsler et al., 2002; Zhao et al., 2003),
@@
    where $\sigma(M)$ is the standard deviation of the linear over-density field on mass scale $M$. Following Chen et al. (2020), we apply Principal Component Analysis (PCA) to $s(z)$ at fixed redshift to reduce dimensionality. The first two principal components (PC1 and PC2) accurately capture both the shape and evolution of a typical MAH, with PC1 representing the overall monotonic growth, while PC2 encodeing a higher-order, non-monotonic deviation. Thus, we use $\dot{M}_{\rm norm}$ to trace recent accretion and (PC1, PC2) to describe the global MAH shape over cosmic time.
  • Halo formation redshift $z_{\rm 1/2}$ is defined as the highest redshift at which the main branch assembled half of its final mass $M_{\rm vir}$ (Wang et al., 2011).
  • Halo 3D shape parameters are derived from the eigenvalues of the inertia tensor (Allgood et al., 2006),
@@
    where the sum runs over all DM particles within the ellipsoid, and $M=\sum_{k}m_{k}$ is the enclosed mass. The eigenvalues of $\mathcal{S}$ correspond to the squared semi-axes ($a\geq b\geq c$), obtained iteratively from the virial sphere until convergence (Tomassetti et al., 2016). We define the 3D axis ratios
@@
  • Environment classification is based on a halo’s location within the cosmic web. We identify web types using the eigenvalues of the deformation tensor (e.g., Hahn et al., 2007; Forero-Romero et al., 2009): the number of eigenvalues exceeding a threshold $\lambda_{\rm th}=0.4$ determines the environment, with 0, 1, 2, and 3 corresponding to voids, sheets, filaments, and knots, respectively. The overdensity field is computed on a $512^{3}$ grid via a cloud-in-cell interpolation, smoothed with a Gaussian filter of $R_{\rm G}=0.5\,h^{-1}{\rm ckpc}$. These parameter choices yield a visually consistent web morphology and agree with previous studies (e.g., Martizzi et al., 2019).
@@ II.3 Galaxy Quantities
II.4 Machine Learning Algorithms
@@
While (semi-)analytical and empirical models have used connections between disk size and halo spin (e.g. Cole et al., 2000; Lacey et al., 2016; Henriques et al., 2020) or concentration (e.g. Nadler et al., 2020; Behroozi et al., 2022; Somerville et al., 2025), it remains an open question which one is more influential (Jiang et al., 2019) and whether incorporating a broader range of halo properties can enhance prediction accuracy (Liang et al., 2025). Moreover, it is not yet known whether disk vertical height and mass fraction, can be robustly inferred from halo properties. Addressing these issues necessitates a systematic investigation of correlations across a large set of halo conditions, a task that can be effectively done by supervised machine-learning (ML) algorithms.
@@
In this study, we use ML algorithms to uncover relations between disk parameters and halo conditions. The targets and features are galaxy quantities and halo quantities, respectively, which are summarized in Table 1. We consider two ML algorithms. We first train the random forest (RF) regressor (Breiman, 2001) for all halo conditions to predict our targets. Then we use SHapley Additive Explanations (SHAP) values (Lundberg and Lee, 2017) to identify important halo properties that affect each prediction. Finally we use symbolic regression (SR, e.g. Koza, 1992) implemented in the package PySR (Cranmer, 2023), with the most important 10 halo properties as identified by the RF regressor as input to derive empirical functions relating the targets and features. The principle of ML algorithms and hyper-parameters that we tune are described in Appendix A. The workflow in this work is presented in Fig. 1.
For training SR models, we adopt a loss function $\mathcal{L}$ that combines a normalized mean squared error term $\mathcal{L}_{\rm NMSE}$ and a normalized quantile loss term $\mathcal{L}_{\rm NQL}$,
$\mathcal{L}=\mathcal{L}_{\rm NMSE}+\mathcal{L}_{\rm NQL}.$ (10)
Here, the normalized mean squared error term takes form
$\mathcal{L}_{\rm NMSE}=(y-\hat{y})^{2}/y^{2},$ (11)
where $y$ is the data and $\hat{y}$ is the predicted values. The normalized quantile loss is given by
$\quad+(1-\tau_{i})\max\Big(-\frac{y-\hat{y}}{y},0\Big)\Bigg],$ (12)
which captures the asymmetric penalties for underestimation and overestimation across multiple quantiles $\tau_{i}$ (0.1, 0.25, 0.5, 0.75, and 0.9). This loss function is able to capture information about the entire conditional distribution of the target variable, rather than only the mean. This is particularly useful when the data contain outliers or are asymmetrically distributed. At early training stage, when the residuals are large, the NMSE term typically dominates and encourages fit of the mean values. As training progresses and residuals become small, the NMSE term decreases, while the NQL term become comparable or more influential, guiding the model towards improved distribution match, avoiding many outliers in the predictions. This loss-function design is adopted in previous astrophysical studies (e.g. Salim et al., 2025), and theoretical considerations can be found in Terven et al. (2025).
Besides the model expressions, PySR also returns the loss $\mathcal{L}$ and complexity $\mathcal{C}$. More complex expressions typically achieve lower loss but are more difficult to implement into semi-analytical models. To balance complexity and loss, we devise a score function to automatically select equations. We first construct normalized loss $\tilde{\mathcal{L}}$ and normalized complexity $\tilde{\mathcal{C}}$ as
$\tilde{\mathcal{L}}=\frac{\mathcal{L}-\min(\mathcal{L})}{\max(\mathcal{L})-\min(\mathcal{L})},\ \tilde{\mathcal{C}}=\frac{\mathcal{C}-\min(\mathcal{C})}{\max(\mathcal{C})-\min(\mathcal{C})},$ (13)
respectively. Then the score is given by
$\text{Score}=\Bigg[w\log(1+\tilde{\mathcal{L}})+(1-w)\log(1+\tilde{\mathcal{C}})\Bigg]^{-1},$ (14)
where $w$ is the weight. This formula first increases with complexity, reaching a maximum, and then decreases with complexity. We choose the equation at the maximum point. We find that the loss decreases very slowly once the model complexity exceeds a certain threshold value. To identify the simplest equation whose complexity is just sufficient to reach this value, we find, by trial and error, a complexity weight of $w=0.7$ is optimal. A higher $w$ results in greater complexity but lower loss at the maximum point.
To avoid overly complicated SR relations and overfitting, we restrict the allowed operators to multiplication ($\times$), plus (++), power ( $\widehat{}$ ), exponential ($\exp$), and base-10 logarithm ($\log$). The maximum expression complexity is set to 100. Further details on the SR parameters are provided in Appendix A.3.
@@
For both RF and SR, we use 80% of the total samples as the training set and the remaining 20% as the test set. All features and target variables are scaled to dimensionless quantities. For the targets such as disk scale radii, disk scale heights, and galaxy sizes, they are normalized by the virial radius, and the dimensionless quantities are referred to as compactness or thickness. The stellar mass is normalized by the virial mass, while the SFR is normalized by the average halo growth rate over Hubble time. For the features, the virial mass is scaled by $10^{10}M_{\odot}$, $M_{\rm vir,10}=M_{\rm vir}/10^{10}{\rm M}_{\odot}$, and the other properties are already dimensionless. Except for disk mass fractions, a base-10 logarithm is applied to the target variables.
@@
III Galaxy-halo connections
In this section, we present the models for predicting galaxy quantities with DM-halo quantities, as identified with the random-forest (RF) and symbolic-regression (SR) algorithms. We first test the performance of the regression models for global properties, specifically, stellar mass, star formation rate (SFR), and size, as a validation of the methods in Section III.1, and then extend the analysis to disk properties in Section III.2. For each relation, we first show the accuracy of our prediction, and then discuss the correlation between the target and the features.
III.1 Predicting galaxy global properties with halo parameters
III.1.1 Stellar mass
@@
III.1.2 Star formation rate
@@ III.1.3 Galaxy size
Feature importance: Unlike stellar mass or SFR, the most important halo property for galaxy compactness is not virial mass but redshift. In addition, the importance of virial mass is much smaller for the hydro halos compared to the DMO halos. For the hydro halos, the role of the Einasto shape index ($\alpha$) is significant, ranking the second, whereas it is not in the top 10 in the DMO case. This difference persists in predictions for disk structures (Section III.2), and implies that $\alpha$ is significantly influenced by the galaxy rather than being a driver of galaxy size. Other commonly important properties across both hydro halos and DMO halos are concentration $c$, spin $\lambda$, and the inner-halo spin $\lambda_{\rm inner}$. A few environmental and assembly-history properties also weakly contribute to the prediction.
@@
III.2 Predicting disk properties with halo properties
III.2.1 Disk size